OpenStudy (anonymous):

Determine the total number of five-card hands that can be drawn from a deck of 52 cards.

OpenStudy (anonymous):

This is a permutations problem. Do you see that if I draw one card there are 52 possibilities?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, after I draw the first card, how many possibilities are left for the second card?

OpenStudy (anonymous):

47

OpenStudy (anonymous):

I took 1 card from 52, how many are left? Why 47?

OpenStudy (anonymous):

53

OpenStudy (anonymous):

Not 53! 51

OpenStudy (anonymous):

51

OpenStudy (anonymous):

Have you ever played cards?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Suppose I give you 5 letters, abcde OK?

OpenStudy (anonymous):

Now, I ask you to pick one. How many possible?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

If we have n objects (cards, letters or whatever) for the first choice there are n possible for the second choice there are n-1 possible and so on

OpenStudy (anonymous):

Back to the cards, for 52 cards, for the first choice there are 52 possible for the second choice there are 51 possible for the third choice there are 50 possible get the pattern

OpenStudy (anonymous):

It's hard if you don't know about cards. So think of the 26 letters of the alphabet, upper and lower case is 52. You draw one, how many left to choose from?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

52

OpenStudy (anonymous):

OK. So in your problem wie have 52 x 51 go ahead..

OpenStudy (anonymous):

2652

OpenStudy (anonymous):

Is that what 52 x 51 x 50 x 49 x 48 equals?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you see why that is the answer?

OpenStudy (anonymous):

311875200

OpenStudy (anonymous):

so it is 2652

OpenStudy (anonymous):

no i think it is 2,598,960

OpenStudy (anonymous):

>>> 52 * 51 * 50 * 49 * 48 311875200

OpenStudy (anonymous):

That's Python. I never calculate by hand if I can help it.

OpenStudy (unklerhaukus):

If there are 52 possible choices of card, and we choose 5 of these The total number of possibilities is \[^{52}C_5=\frac{52!}{5!(52-5)!}=\frac{52!}{5!47!}=\frac{52\times51\times50\times49\times48}{5\times4\times3\times2\times1}=\frac{311875200}{120}=\] their are 52! possible ways to draw all the cards the cards that have not been drawn are (52-5)! we divide by the number of choices because this is a combination problem not a permutation problem ~ the order does not matter

OpenStudy (anonymous):

Yep, order doesn't matter. My mistake.